The angle theta is formed by the positive x axis and a ray through the point (5, -12). Find the value of sin(theta), cos(theta), and tan(theta)
Answer by mathstutor494(120)

Ans :
sin(theta)= -5/13
cos(theta)= 12/13
tan(theta)= -5/12
If you need more help, email me at mathstutor494@gmail.com